OpenStudy (prettygirl_shynice):

5gal 2qt + 4 gal 5qt (there are 4 quarts in one gallon) but explain how i do it

OpenStudy (prettygirl_shynice):

@pooja195 ,@jigglypuff314

OpenStudy (prettygirl_shynice):

@jigglypuff314

jigglypuff314 (jigglypuff314):

have you learned how to add mixed fractions before? :) like 2 1/3 + 1 2/3 = 3 3/3 = 4 ?

OpenStudy (prettygirl_shynice):

yea it is eazy

OpenStudy (prettygirl_shynice):

@jigglypuff314 yea it is eazy

jigglypuff314 (jigglypuff314):

this is the same idea since there are 4 quarts in a gallon the fractions would be over 4 so "5gal 2qt + 4gal 5qt" = 5 2/4 + 4 5/4 = ?

OpenStudy (prettygirl_shynice):

that 10 3/4

OpenStudy (prettygirl_shynice):

so its 10 gallos 3 quarts

jigglypuff314 (jigglypuff314):

perfect ^_^

OpenStudy (prettygirl_shynice):

thx

OpenStudy (prettygirl_shynice):

need help with one more